### v3

# VilBias: A Study of Bias Detection through Linguistic and Visual Cues , presenting Annotation Strategies, Evaluation, and Key Challenges

Abstract
The integration of Large Language Models (LLMs) and Vision-Language Models (VLMs) opens new avenues for addressing complex challenges in multimodal content analysis, particularly in biased news detection. This study introduces ViLBias, a framework that leverages state-of-the-art LLMs and VLMs to detect linguistic and visual biases in news content. We present a multimodal dataset comprising textual content and corresponding images from diverse news sources. We propose a hybrid annotation framework that combines LLM-based annotations with human review to ensure high-quality labeling while reducing costs and enhancing scalability. Our evaluation compares the performance of state-of-the-art SLMs and LLMs for both modalities (text and images) and the results reveal that while SLMs are computationally efficient, LLMs demonstrate superior accuracy in identifying subtle framing and text-visual inconsistencies. Furthermore, empirical analysis shows that incorporating visual cues alongside textual data improves bias detection accuracy by 3–5%. This study provides a comprehensive exploration of LLMs, SLMs, and VLMs as tools for detecting multimodal biases in news content and highlights their respective strengths, limitations, and potential for future applications.
keywords:
multimodal bias detection , large language models , vision-language models ,media bias , hybrid annotation
1 Introduction
The rapid development of Large Language Models (LLMs), such as the OpenAI GPT series and Meta LLaMA, has profoundly influenced natural language processing (NLP) [1]. These models demonstrate remarkable capabilities in generating coherent text and understanding complex language patterns. More recently, researchers have begun integrating LLMs with additional modalities, such as images, audio, and video, which leads to the emergence of multimodal LLMs [2]. These systems excel at tasks requiring joint inference from textual and visual inputs, including image captioning, visual question answering, and content moderation [3].
In the realm of news media, detecting bias remains a pressing concern [4]. Media outlets frequently employ subtle linguistic framing and selective visuals to shape public perception [5]. While traditional text-based methods for bias detection identify patterns in phrasing and emphasis, they often overlook the complex interplay between textual and visual elements. In contrast, multimodal LLMs [6] can analyze multiple modalities simultaneously, offering a more holistic understanding of the data. For bias detection, these systems examine how written narratives interact with accompanying visuals to frame events, highlight specific viewpoints, or influence the dissemination of information. A working example is shown in Figure 1.
Although multimodal fake news detection has received significant attention [7, 8], bias detection remains underexplored; likely due to its subjective nature of the topic, which highlights the need for targeted research in this area. In this work, bias refers to any systematic partiality or distortion in textual or visual content that shapes how an event, individual, or group is portrayed. Such partiality manifests when one perspective is consistently favored, while other viewpoints are undermined, trivialized, or dismissed. This notion of bias is informed by multiple theoretical underpinnings in communication, psychology, and media studies, such as framing theory [9], social identity theory [10], and recent developments in computational science [11]. In multimodal contexts, these biases can be more subtle that are seen not just in written language (e.g., word choice, framing of facts) but also through images, videos, and other visual cues (e.g., cropping, staging, or emotive imagery). Consequently, identifying and categorizing such biases requires a nuanced approach that integrates both theoretical perspectives and data-driven techniques.
Figure 1: Examples of biases in news. Left: Combining image and text to amplify fear—an illustration of multimodal storytelling in conspiratorial narratives. Right: The role of visual and textual multimodality in propagating bias through a fabricated claim supported by engineered imagery. Although the text alone may not exhibit bias, the image demonstrates how multimodality can reveal potential biases and manipulative framing.
A main obstacle in multimodal bias detection is the lack of high-quality, labeled datasets for training and evaluation [12]. Expert annotations, while considered the gold standard, are both costly and time-consuming to produce. Although crowdsourcing can offer efficiency at scale, it often fails to capture the subtlety required for bias detection [13], leading to noisier datasets. Overcoming this limitation is vital for improving the accuracy and reliability of multimodal bias detection systems. The goal of this study is therefore two-fold:
  1. 1.
    To present our hybrid AI-human annotation methodology for creating a scalable, high-quality multimodal bias dataset.
  2. 2.
    To evaluate various LLMs and Vision-Language Models (VLMs) in identifying subtle biases across textual and visual content.
We introduce ViLBias (VIsual and Linguistic cues for Bias detection), a framework designed for both labeling and evaluating LLMs and VLMs on nuanced bias detection. First, we propose a hybrid annotation strategy combining an AI-based labeling process with human oversight to ensure high-quality, multimodal annotations. We then evaluate state-of-the-art (SoTA) LLMs and VLMs in performing binary bias classification on our dataset. This framework builds on our released multimodal dataset, which pairs textual content with visuals from diverse news sources. The goal is to improve label quality, address nuanced biases, and reduce annotation costs .
This study addresses the following research questions:
  1. 1.
    Does multimodality add value compared to a single modality, such as text alone?
  2. 2.
    Are Small Language Models (SLMs) [^1] The definition of SLMs varies in the literature [14]; for our purposes, SLMs are models predating the ChatGPT era that rely on an encoder-only architecture. more effective as bias classifiers, or do more recent LLMs and VLMs outperform them?
Contributions
  • In this study, we present the design and implementation of a hybrid annotation approach that integrates human expertise with LLM-based annotations. We leverage open-source LLMs as annotators and outline a structured methodology for their effective use in annotation tasks, an area that remains relatively unexplored. While some of the images and text samples are drawn from our prior work [15], this study introduces a distinct annotation framework, employing multiple LLMs to improve reliability and robustness. To support reproducibility, all related data and code are available [^2] The code is accessible at https://github.com/shainarazavi/VILBias. Data can be shared upon reasonable request.
  • We assess the performance of SoTA LLMs and VLMs in detecting biases within this multimodal dataset, comparing encoder-only SLMs with autoregressive LLMs for a bias-classification task. We also explore differences in model performance across both modalities (text and images) to identify optimal usage scenarios.
Our experiments show that incorporating visual cues alongside text improves bias detection performance by 3–5%. While SLMs perform competitively in classification tasks, LLMs demonstrate stronger generative capabilities, particularly when working with limited data. These findings underscore both the promise and the current limitations of multimodal bias detection approaches.
The remainder of this paper is organized as follows: Section 2 discusses related work in multimodal bias detection. Section 3 details our hybrid annotation methodology and the ViLBias framework. Section 4 presents experimental results, and Section LABEL:sec:conclusion concludes with insights and future directions.
2 Related Work
The study of fake news can be categorized into disinformation, misinformation [16], and, less frequently, bias news [17]. A clear distinction between biased and fake news is essential for effective media analysis. Biased news typically presents information with a particular slant, emphasizing certain viewpoints while downplaying others, whereas fake news consists of entirely fabricated content intended to mislead readers. This work primarily focuses on bias detection in news.
Research in this area often employs SLMs and LLMs, treating it as a classification task. SLMs are compact, task-specific models with fewer than 100 million parameters, whereas LLMs are larger and more versatile, typically comprising billions of parameters [1]. While there is no universal consensus on the exact distinction between SLMs and LLMs, as noted in [14], this study adopts a practical categorization: models developed before ChatGPT are referred to as SLMs, while those introduced post-ChatGPT are classified as LLMs.
Annotation Process
A significant challenge in this area of research is the scarcity of labeled data. Some studies address this issue by utilizing human annotations, often sourced from crowd workers [18], while others rely on domain experts [19] or, in rare cases, AI-generated labels [20]. The process of manual labeling typically involves human annotators with extensive knowledge, making it costly, time-consuming, and prone to errors. To mitigate these challenges, research community start exploring AI based methods that can help scale the labeling process. For example, the use of openAI GPT is seem lately for annotation of the data [20]. The goal of this work is to alleviate data scarcity in specific domains [12].
The integration of AI in data labeling processes introduces both opportunities and challenges that must be navigated with care [21]. Human oversight remains essential to ensure the quality and fairness of labeled data, as expert annotators play a crucial role in reviewing and validating AI-generated annotations. Some recent works that consider AI and human based annotations include [22, 23, 24]. Studies such as [25] and [26] relied on AI-only annotations. Hybrid SLM-LLM approaches, as seen in [27] and [28, 29], used human + AI annotations.
Table 1: LLM-based annotations in related work.
Paper Labelling Method LLM Used Problem Solved Evaluation Metric
[30] Prompts OpenAI models General, classification, text generation Human verification
[31] Prompts GPT-3.5, GPT-4 mainly, open-sourced did not work Political claim detection Accuracy, Cohen’s Kappa
[32] Prompts, explain then annotate GPT-3.5 Question answering, crowd-sourced data Accuracy, recall, MRR, Likert scale
[20] Prompts ChatGPT Classification, comparison with MTurk workers Accuracy, intercoder agreement
[21] Prompts GPT-3.5, GPT-4, LLaMA 1 Classification, comparison with MTurk workers Accuracy, intercoder agreement
[33] Prompts GPT-3.5, GPT-4 Bias detection in hate speech classification Match/mismatch with ground truth
[34] Prompts, automatic reweighting GPT, GPT-3, GPT-4 Efficiency and reliability in annotation Precision, recall, F1, std deviation
[35] Prompts GPT-4 Image captioning, text style transfer BLEU, ROUGE, BERTScore, others
[36] Prompt, weak supervision T0++, GPT-3 Text classification, spam detection Precision, recall
[37] Prompts GPT-3, InstructGPT Token classification Macro F1, accuracy
[38] Prompts GPT-3 Comparison with human annotations on NLG, NLU tasks Accuracy, ROUGE
[39] Prompts GPT-3 Comparison with human annotations on SST-2 Precision, recall, F1, time, cost
[40] Prompts GPT-3.5-turbo Translation BLEU, ROUGE
Table 2: Summary of Multimodal Models for Fake/Biased News Detection
Model Method Summary Free Dataset [^3] COCO, etc., refer to notable datasets due to space constraints.
CLIP [41] CNN, ResNet50 Predicts captions by selecting the most similar candidate based on similarity scores. Yes aYahoo, ImageNet, SUN
MiniGPT4 [42] Vicuna Describes an image based on an input image and question. Yes RefCOCO
Fuyu-8b [43] Decoder-trans Generates captions by processing images and text-based questions. Yes COCO, etc.
imageBind-LLM [44] LLaMA Handles multiple modalities to generate image descriptions. Yes COCO, etc.
MultiModal-GPT [45] CLIP and LLaMA Creates image captions and answers image-based queries. Yes COCO, etc.
Kosmos-2 [46] Method Accepts images and questions to describe objects. RefCOCO, etc.
NExT-GPT [47] ImageBlind Works with multiple modality combinations. No MosIT Data
DreamLLM [48] CLIP Capable of generating free-form content. No COCO, etc.
SPHINX [49] LLaMA 2 Provides effective image descriptions. Yes LAION-400M
BLIP-2 [50] BLIP An enhanced version of BLIP for image descriptions. Yes COCO
LLaVA [51] Vicuna Processes image and language instructions for image-based answers. Yes COCO
FakeNewsGPT4 [52] LVLM Detects news manipulation using specialized LVLM knowledge. No DGM
LEMMA [53] LVLM Enhances LVLM with external knowledge for misinformation detection. No Twitter, Fakeddit
MCNN [54] BERT, ResNet50 Combines BERT and ResNet for feature extraction to detect fake news. No D1-D4
HAMMER [55] Attn, FFN Focuses on modality interactions for misinformation detection. Yes DGM
LLMs as Annotators
Comparative analyses of multiple studies indicate that while LLMs provide cost- and time-saving benefits for classification tasks, they also encounter challenges such as lack of representativeness, bias, sensitivity to prompt variations, and a strong preference for the English language [56]. Evaluations comparing LLM annotations with those of human experts and novices reveal moderately high agreement with human experts across most quality dimensions. Additionally, incorporating LLMs as supplementary annotators can enhance inter-annotator agreement, underscoring their value in automated quality assessment tasks [13]. Frameworks that integrate LLMs into active learning loops have also been proposed to improve annotation efficiency [57]. By selecting the most informative samples for LLM annotation and employing example reweighing techniques, these frameworks enable task-specific models trained on LLM-generated labels to outperform original models, even with minimal annotated examples.
The performance of multilingual LLMs in annotating low-resource language content has been evaluated for tasks like natural language inference, sentiment analysis, and news classification. Results indicate substantial variability, with off-the-shelf multilingual LLMs that shows poor performance in certain tasks, particularly in low-resource settings [58]. The feasibility of employing LLMs as replacements for human annotators has also been investigated. Zero-shot performance across various tasks reveals that models trained with human-generated labels generally outperform those trained with LLM-generated labels. Recent work on fake news and bias detection using LLMs for annotations or evaluations is summarized in Table 1. As highlighted in the table, these annotation methods primarily focus on text-based tasks. To the best of our knowledge, this study is the first to extend such approaches to a multimodal context.
Multi-modal Bias Detection
The integration of multiple modalities in media content necessitates advanced methods for detecting bias. Research on multimodal AI emphasize the need for datasets and models that effectively address biases arising from the combination of different data types [59, 60]. SoTA approaches include CLIP-guided learning techniques for multimodal fake news detection, which leverage cross-modal capabilities to enhance accuracy by analyzing both textual and visual elements [61]. Other related works capture fine-grained connections between images and text through multimodal interaction and graph contrastive learning [62]. A detailed literature review of techniques for detecting multimodal fake and biased news, noting the relatively limited research on the latter, is summarized in Table 2.
3 Multi-Stage Annotation Methodology
3.1 Problem Definition
In this work, we focus on the task of multimodal bias detection in news media. Formally, let $\mathcal{T}$ represent the textual component (e.g., an article or caption) and $\mathcal{V}$ represent the visual component (e.g., an accompanying image). We define the problem as follows:
Input:
A pair $(\mathcal{T},\mathcal{V})$, where $\mathcal{T}$ is a textual snippet and $\mathcal{V}$ is a corresponding visual element.
Output:
A binary label $y\in\{0,1\}$ indicating the presence (1) or absence (0) of bias in the combined textual and visual content.
Here, bias is defined as a systematic partiality or distortion in how an event, individual, or group is portrayed. Such partiality may manifest when one perspective is consistently favored, while other perspectives are undermined, trivialized, or dismissed. Our focus on multimodality recognizes that biases often emerge from the interplay between textual framing (e.g., word choice, omission of details) and visual cues (e.g., cropping, emotive imagery).
Our framework, ViLBias, is depicted in Figure 2, which illustrates three main stages:
  1. 1.
    Data Sourcing and Preprocessing
  2. 2.
    Annotations via LLM/VLM and Human Validation
  3. 3.
    Evaluations
Figure 2: ViLBias Framework. The pipeline comprises three main stages: (1) Data Collection and Preprocessing, (2) Automated and Human-Annotated Labeling, and (3) Final Evaluation.
3.2 Data Compilation
We selected a range of reputable and widely recognized news organizations for data collection, guided by prior research on media bias and source credibility [5, 63]. Our chosen outlets, CNN, BBC, The New York Times, and Al Jazeera, cover a broad spectrum of geographic regions, editorial policies, and audience trust profiles. Appendix LABEL:tab:news_sources provides detailed information about these sources.
Data were collected between May 6, 2023, and September 6, 2024, capturing evolving patterns of disinformation under shifting global events. To ensure compliant and ethical data practices, we adhered to privacy regulations, complied with website terms of service, and followed established guidelines [64].
Preprocessing focused on enhancing dataset quality and relevance. We removed duplicate articles to eliminate redundancy and filtered out multimedia-based items (e.g., URLs ending in /video, /slide, or /gallery) to concentrate on text-rich content. Articles with fewer than twenty sentences were also discarded to ensure sufficient textual depth for meaningful analysis. Each retained article was paired with its first encountered image, providing a consistent multimodal reference point.
Some of the images and text samples in this study are sourced from our prior work in same project [15]. However, the labeling scheme used here differs, as it incorporates multiple LLMs for annotation, enhancing robustness and reliability.
3.3 Multiple Annotations
To efficiently annotate a large, complex dataset, we integrated automated annotators in two categories:
  1. 1.
    LLMs for textual annotations, and
  2. 2.
    VLMs for multimodal (text-image) labels.
The prompt templates used for text-based annotations are provided in Appendix B.
A key feature of our approach is a two-step voting mechanism that increases annotation reliability. First, each LLM was queried three times per item to address stochastic variation inherent in LLMs [65], and the annotation from each LLM was determined by majority vote. Next, using multiple LLMs (three, in our case), the final annotation was decided by a second majority vote across these models.
Formally, let $A_{m}$ be the set of annotations from model $m$ after three queries. The final annotation $A_{\text{final}}$ is:
$A_{\text{final}}=\arg\max_{a\in A}\sum_{m\in M}\mathbb{1}(A_{m}=a),$
where $M$ is the set of LLMs (an odd number), $A$ is the set of all possible annotations, and $A_{m}$ is determined as:
$A_{m}=\arg\max_{a\in A}\sum_{i=1}^{3}\mathbb{1}(A_{m}^{(i)}=a),$
with $A_{m}^{(i)}$ being the $i$-th query response from model $m$. The indicator function $\mathbb{1}(A_{m}^{(i)}=a)$ equals 1 if the annotation matches $a$, and 0 otherwise. This voting approach mitigates stochastic variability within and across models, ensuring robust final annotation decisions.
Human Validation
Despite automation, human oversight remains essential for validating model outputs and calibrating the annotation pipeline. We enlisted 12 reviewers, comprising three undergraduates, four master’s students, two Ph.D. holders, and three senior researchers across computer science, media studies, political science, and linguistics. Two subject-matter experts led this process, guided by bias identification protocols outlined in Appendix B. These guidelines ensured that reviewers evaluated both linguistic and visual biases as defined by the U.S. Equal Employment Opportunity Commission [66].
Reviewers assessed a stratified sample of 10,000 annotations, classifying them as Acceptable, Needs Improvement, or Incorrect, and provided written rationales where necessary. This feedback informed iterative improvements to model prompts and annotation strategies. Disagreements were resolved through consensus discussions. After refining our approach, the inter-annotator agreement (IAA), calculated using Cohen’s Kappa, between automated labels and expert judgments reached 0.72, indicating substantial alignment.
Final Dataset
The resulting dataset is organized to support bias analyses across both textual and multimodal dimensions. Each article is assigned a stable, non-identifying unique identifier derived from the first ten digits of the SHA256 hash of its headline. Core fields include:
  • outlet, headline, article text (snippet), image, URL
To capture image semantics, we use a VLM (BLIP [67]) to generate a textual image description. The final dataset contains approximately 41,412 articles, each paired with its primary image. Two key annotation fields—nlp label and nlp-image label—indicate whether the text or the combined text-image pair is likely to exhibit bias. This schema enables diverse downstream tasks, from bias detection to explanatory modeling.
Table 3 summarizes key statistics and observed patterns. Figure 3 provides sample articles and their associated images, illustrating the multimodal scope of this collection.
Table 3: Data Analysis Summary.
Metric Value
Number of Rows 40,945
Number of Columns 7
Field Names unique_id, headline, url, article_text, image, nlp_label, nlp-image_label
Top Outlets (by count) Financial Times (4895), USA TODAY (3320), CNN (3067), The Guardian (2833), Forbes (2482)
Unique Values in nlp_label 2
Distribution of nlp_label {‘Bias’: 28,281, UnBias’: 12,664}
Unique Values in nlp-image_label 2
Distribution of nlp-image_label {‘Bias’: 22,974, ‘UnBias: 17,971}
Unique Images 40,945
Figure 3: Sample headlines and their associated images. Each image corresponds to a specific news article, illustrating the multimodal structure of the dataset.
Annotation Quality
We assess the annotation quality to address the research question: How accurate are annotations generated by LLM-based systems? To evaluate this, we selected a representative sample of 1000 records from our labeled dataset. This sample size was chosen to provide a statistically significant but manageable dataset for analysis and represents a balanced distribution between instances labeled to be biased or unbiased. The selected records were independently labeled by our team, who had no prior knowledge of the existing annotations. To assess the quality of LLM-generated annotations, we calculated precision, recall, F1 score, and accuracy, metrics widely used for evaluating classification performance.
Table 4: Performance metrics based on a majority vote from the dataset and human annotated sample, evaluated using a sample of 1000 examples. Metrics include precision, recall, F1 score, and accuracy for both text and multimodal labels.
Metric Text Label (%) Multimodal Label (%)
Precision 85.0 80.0
Recall 88.0 83.0
F1 Score 86.5 81.5
Accuracy 87.0 84.0
The results presented in Table 4 showcase the strong performance of LLM-based annotations, with all metrics surpassing 80%. This underscores the effectiveness of LLMs in handling both text-based and multimodal annotation tasks. Interestingly, text labels outperformed multimodal labels, achieving higher precision (85.0% vs. 80.0%), recall (88.0% vs. 83.0%), F1 score (86.5% vs. 81.5%), and accuracy (87.0% vs. 84.0%). These results highlight that multimodal labeling poses additional challenges, likely due to the complexities of integrating textual and visual information. Nevertheless, multimodal annotations remain competitive, with an accuracy of 84.0%, demonstrating their potential to capture subtle biases and richer context across modalities. The small gap between precision and recall for both text and multimodal labels reflects balanced and consistent performance, further validating the reliability of these annotations. Overall, these findings underscore the promise of multimodal models while also emphasizing the need for continued research to improve their ability to effectively leverage interactions between text and visuals.
4 Experimental Setup
4.1 Setting
For this work, we conducted our experiments using computing setup equipped with NVIDIA A100 Tensor Core GPUs to leverage efficient parallel processing capabilities for training and inference. The software environment included Python 3.10 with essential deep learning libraries such as PyTorch 2.0, Transformers by Hugging Face for model implementation, and OpenCV for image preprocessing. The models were trained and evaluated on Ubuntu 22.04 LTS running CUDA 12.1 for optimal GPU acceleration.
4.2 LLMs as Annotators
For text-only tasks, we utilized open source models such as Llama-3-8B-Instruct, Llama-3.1-8B-Instruct [68], and Mistral-7B-Instruct-v0.3 [69]. For text-image tasks, our approach integrated LLaVa-NeXT [70], Phi-3-Vision-128k-Instruct [71], and MiniCPM-V-2_6 [72]. To handle potential conflicts, we ensured an odd number of LLMs were used as annotators, enabling effective resolution of ties. This means that when there are disagreements among the annotators, the majority vote determines the final annotation, ensuring a definitive outcome without ambiguity. Our prompts for annotation are given in Appendix B.
4.3 Baseline Models and Evaluation Metrics
To conduct a comprehensive performance evaluation, we categorized models into SLMs and LLMs. In our study, SLMs encompass transformer architectures such as BERT [73], RoBERTa [74], DistilBERT [75], GPT-2 [76], and BART [77]. We also included multimodal configurations such as XLNet paired with VGG-19, BERT with ResNet-34, and BERT integrated with CLIP.
LLMs in our framework are represented by text-only models including LLama3.2-1B (both instruct and base variants) [78], Mistral-v0.3-instruct 7B [69], and Qwen2.5-7B-Instruct [79]. Furthermore, we examined VLMs such as InstructBlipV [80], Phi-3-vision-128k-instruct [71], Pixtral-12B-2409 [81], LLaVA-1.6 [82], and Llama-3.2-11B-Vision-Instruct [78]. This selection spans from compact, task-specific designs to versatile, billion-parameter frameworks, encompassing both text-only and multimodal platforms.
For evaluation, we tested each model under zero-shot and few-shot (5-shot) scenarios , and we also fine-tuned them using our entire dataset. We utilized an 80/10/10 split for training, validation, and testing subsets, respectively, ensuring a balanced representation of classes in each. Additionally, we implemented 5-fold cross-validation to verify the models’ robustness and employed stratified sampling to maintain uniform class distribution across the folds. Our metrics for performance assessment included standard ML metrics, such as precision, recall, F1-score, accuracy, and some qualitative evaluation through human validation.
4.4 Hyperparameters
All models were implemented using PyTorch, leveraging specific configurations to optimize performance and resource utilization. Key configurations include:
  • Data Type: torch.bfloat16 was used to enable mixed-precision training, striking a balance between memory efficiency and numerical accuracy.
  • Efficient Attention Mechanism: flash attention 2 was employed (where applicable) to accelerate attention computations while reducing memory overhead, particularly for larger models.
  • Device Mapping: Models were mapped to cuda-enabled devices to utilize GPU acceleration for faster computations.
Additionally, we employed quantization techniques when fine-tuning LLMs and VLMs to reduce memory requirements and speed up inference, particularly for large-scale models. For SLMs, regular fine-tuning without quantization was applied, as these models have a smaller parameter space and are computationally less demanding.
Table 5: Key Parameters for Training, Response Generation, and Evaluation
Parameter Details
Generation Temperature: 0.2–0.4, Maximum token limit: 1024
Batch Sizes Prompt Evaluation: 100
BERT Training: 32 (15 epochs, learning rate: $2\times 10^{-5}$, weight decay: 0.01, early stopping patience: 5)
Fine-Tuning: 1 (3 epochs, learning rate: $2\times 10^{-4}$, gradient accumulation: 1, max gradient norm: 0.3)
Learning Rate Schedulers Constant with warmup ratio of 0.03
Precision Mixed precision enabled (bf16=True)
Loss Functions CrossEntropyLoss (default)
Focal Loss (alpha=1, gamma=2) for specific tasks
Evaluation Metrics Accuracy, Precision, Recall, F1 Score, Confusion Matrices (Batch Size: 32)
Quantization 4-bit precision, double quantization enabled, nf4 quantization type
Environment and Reproducibility CUDA-enabled devices, CPU fallback when necessary
Random seed set to 0
Parallel tokenization disabled (os.environ["TOKENIZERS_PARALLELISM"] = "false")
5 Results
5.1 Overall Performance Evaluation
In this experiment, we investigate the comparative classification effectiveness of SLMs versus LLMs and VLMs. Our study specifically examines model performance across different configurations and modalities. We evaluated the models in two primary settings:
  1. 1.
    Text-only classification: Using small and large language models.
  2. 2.
    Multimodal classification: Employing small and large vision-language models.
Model Config Precision Recall F1 Score Accuracy
Small Language Models (Text-Only)
BERT-base-uncased FT 87.10% 86.50% 86.80% 86.90%
DistilBERT FT 84.30% 83.40% 83.85% 85.00%
RoBERTa-base FT 89.20% 87.90% 88.54% 89.50%
GPT-2 FT 85.50% 85.00% 85.24% 85.10%
BART FT 85.80% 85.70% 85.75% 85.80%
Large Language Models (Text-Only)
LLama3.2-1B-Instruct 0-shot 72.50% 73.40% 72.95% 67.80%
5-shot 76.20% 74.00% 75.08% 71.10%
IFT 84.00% 82.30% 83.14% 82.50%
Mistral-v0.3 7B-Instruct 0-shot 68.00% 65.00% 66.47% 67.50%
5-shot 80.10% 78.90% 79.49% 77.80%
IFT 86.20% 83.40% 84.78% 80.90%
Small Vision-Language Models (Text + Image)
XLNet + VGG-19 FT 72.00% 68.20% 70.05% 77.10%
BERT + ResNet-34 FT 75.80% 71.50% 73.59% 79.40%
BERT + CLIP FT 81.30% 73.40% 77.15% 84.20%
DistilBERT + CLIP FT 68.50% 64.00% 66.18% 74.90%
RoBERTa-base + CLIP FT 84.50% 81.40% 82.90% 83.60%
Large Vision-Language Models (Text + Image)
Phi-3-Vision-128k-Instruct 0-shot 70.40% 66.00% 68.12% 69.80%
5-shot 73.20% 71.00% 72.08% 70.50%
IFT 76.80% 78.10% 77.44% 74.00%
LLaVA-v1.6-Mistral-7B 0-shot 62.50% 60.80% 61.64% 62.70%
5-shot 68.10% 67.00% 67.54% 65.20%
IFT 75.40% 74.60% 75.00% 76.10%
LLama3.2-11B-Vision-Instruct 0-shot 65.00% 68.90% 66.88% 68.40%
5-shot 73.40% 74.20% 73.80% 72.10%
IFT 80.20% 81.90% 81.04% 82.50%
Table 6: Performance of small and large language models (text-only) and vision-language models (text + image) across configurations: 0-shot, 5-shot, fine-tuning (FT), and instruction fine-tuning (IFT). Metrics are presented in percentages.
Our evaluation reveals several key insights:
Text-only models: Table 6 provides an overview of how both SLMs and LLMs perform on text classification tasks. Among the SLMs, RoBERTa-base emerged as the best-performing model after fine-tuning, achieving an F1 score of 88.54% and an accuracy of 89.50%. Other SLMs, such as BERT and DistilBERT, also achieved strong results, although their simpler architectures and fewer parameters led to slightly lower performance.
For LLMs, Llama3.2 demonstrated robust performance, closely followed by Mistral. Their results underline the importance of training methodology and data size: instruction fine-tuning (IFT) outperformed both 5-shot and 0-shot approaches, showing that larger data samples and more demonstrations improve LLM training outcomes.
Although LLMs rely on patterns in training data to predict outputs, their predictions are not inherently probabilistic in the sense of randomness [83, 65]. During tasks like text classification, both SLMs and LLMs typically select the most likely output (argmax of the predicted distribution), which makes their behavior deterministic given fixed inputs and parameters. However, in tasks like text generation, probabilistic decoding methods (e.g., sampling with temperature or top-k sampling) may be used to introduce variability. Achieving a 100% success rate remains unrealistic for both SLMs and LLMs due to factors like limited training data, inherent ambiguities in language, and the complexity of interactions among parameters. While SLMs excel in specialized tasks due to focused optimization, LLMs rely on diverse and extensive training data to fully leverage their broader capabilities. These factors contribute to the modest performance gap observed between SLMs and LLMs for certain tasks.
Overall, these findings underscore that no single model or training approach is universally optimal. Instead, performance depends on a combination of factors, including the model’s architecture, the size and diversity of the training data, and the specifics of the fine-tuning process.
Multimodal Models: Table 6 presents the comparative performance of SLMs and large VLMs for multimodal classification tasks. Among small models, the combination of BERT with CLIP achieved the highest F1 score of 77.15% and an accuracy of 84.20% under the fine-tuning (FT) configuration. This result highlights the effectiveness of combining strong text encoders with vision models for multimodal tasks.
Large models demonstrated even greater potential, particularly with IFT. LLama3.2-11B-Vision-Instruct emerged as the best-performing model in this category, achieving an F1 score of 81.04% and an accuracy of 82.50%. The improvements underscore the importance of IFT in aligning multimodal capabilities, as it consistently outperformed both 0-shot and 5-shot configurations across models.
The results also reflect performance variations across configurations—0-shot, 5-shot, fine-tuning (FT), and IFT. For instance, small vision-language models like XLNet combined with VGG-19 showed relatively modest results, with an F1 score of 70.05% and accuracy of 77.10%. In contrast, larger models such as Phi-3-Vision-128k-Instruct demonstrated substantial improvements with IFT, reaching an F1 score of 77.44% and an accuracy of 74.00%. Similarly, LLaVA-v1.6-Mistral-7B showed the greatest gain when transitioning from 0-shot to IFT, reinforcing the value of tailored fine-tuning approaches for optimizing performance.
These results highlight the trade-offs between smaller and larger models, where small models often perform better in efficiency and simplicity, while large models leverage their scale and advanced training methodologies to achieve superior performance.
5.2 Ablation Studies
To better understand which components and training configurations most significantly impact model performance, we conducted a series of ablation experiments on fine-tuned both text-only and multimodal classification tasks. Specifically, we focused on three main aspects: data size, fine-tuning techniques, and multimodal fusion strategies.
5.2.1 Ablation 1: Training Data Size
We varied the size of the training dataset for a representative LLM (LLama3.2-1B-Instruct) and a SLM (RoBERTa-base) to examine how data quantity influences performance. The results in Figure LABEL:fig:ablation1 show that LLama3.2-1B outperforms RoBERTa-base in both F1 score and accuracy across all training data fractions. Additionally, LLama3.2-1B maintains better stability as the training data is reduced, indicating its robustness in low-resource scenarios, whereas RoBERTa-base exhibits a steeper decline in performance with reduced data availability. This shows that LLMs are more stable, and instruction fine-tuning further enhances their performance and resilience to limited training data.
5.2.2 Ablation 2: Fine-Tuning Techniques (Parameter-Efficient Methods)
Fine-tuning LLMs such as LLama3.2-1B and multimodal models like LLama3.2-11B-vision-Instruct can be computationally expensive due to the need to update all parameters during training. To address this, parameter-efficient fine-tuning (PEFT) [84] methods like Low-Rank Adaptation (LoRA) and Prefix-Tuning aim to reduce computational overhead by updating only a small subset of model parameters.
Full fine-tuning (FT) updates all model parameters during training, requiring high computational and memory resources but generally achieving the best performance. LoRA introduces low-rank trainable layers to the frozen pretrained weights, allowing significant parameter savings while maintaining performance. Prefix-Tuning appends trainable prefix embeddings to the input representations, leaving the main model frozen and thus reducing computational demands.
For this experiment, we used LLama3.2-1B-Instruct for text-only tasks and LLama3.2-11B-vision-Instruct for multimodal tasks (text + image). The models were trained on 80% of the dataset, with 20% held out for evaluation. The evaluation metrics used are pecision, recall, F1 score, and accuracy. The goal was to evaluate if these parameter-efficient methods could achieve competitive performance compared to full fine-tuning.
Table 7: Performance Comparison of Fine-Tuning Techniques for LLaMA 3.2 Models.
Model Method F1 Score Accuracy Drop (%)
LLama3.2-1B (Text) Full FT 83.1% 82.5% -
LoRA 82.3% 81.8% 0.8%
Prefix-Tuning 80.7% 80.5% 2.4%
LLama3.2-11B (Text + Vision) Full FT 81.7% 83.0% -
LoRA 81.0% 82.0% 0.7%
Prefix-Tuning 80.2% 81.2% 1.5%
Table 7 shows that full fine-tuning delivers the best performance across both text-only and multimodal tasks but comes at the cost of higher memory and computational requirements. LoRA achieves competitive performance with only a 0.7-0.8% drop in F1 score and accuracy, making it a highly practical alternative for resource-constrained environments. Prefix-tuning shows slightly more degradation, with a 1.5-2.4% drop in F1 score and accuracy, but remains viable when computational resources are extremely limited. These results demonstrate that PEFT methods can achieve near-parity with full fine-tuning while significantly reducing the computational overhead, making them suitable for deployment on edge devices or in scenarios with limited training infrastructure.
5.3 Qualitative Analysis and Bias Identification
In this section, we analyze potential biases in model predictions by examining how these models reason about certain statements. We elicit the chain-of-thought from LLMs and VLMs to determine whether their reasoning (1) relies on stereotypes or over-generalizations, or (2) correctly identifies or dismisses biases where relevant. We use our fine-tuned Llama3.2 ($1$B and $11$B), Mistral-v0.3-7B, and Phi3-vision model for generating predictions.
5.3.1 Text-based Qualitative Analysis and Bias Identification
Table 8 summarizes typical examples of each model’s responses. For each input statement, the model predicts whether there is Likely or Unlikely bias, then provides a brief rationale. We focus on two main aspects:
  1. 1.
    Misclassified Bias: Cases where the model either overestimates or underestimates the presence of bias due to stereotype-based reasoning or a failure to recognize marginalized perspectives.
  2. 2.
    Correct Classification: Cases where the model appropriately recognizes or dismisses bias based on the content and context of the statement.
Model Predic Example and Reasoning Bias Observed
Llama3.2-1B Bias Input: “The nurse provided a groundbreaking solution during a critical surgical operation.”
Reasoning (model): “Nurses rarely contribute to surgical innovations; doctors usually lead.”
Issue: Stereotype-based reasoning. The model assumes nurses cannot be innovators, reinforcing traditional roles and overlooking individual competence.
Llama3.2-1B Unbias Input: “The committee led by women executives proposed significant policy reforms.”
Reasoning (model): “Focus on reforms; no inherent bias.”
Issue: Fails to acknowledge or explore the possibility of gender-based biases in leadership perceptions; too dismissive.
Llama3.2-11B Unbias Input: “The nurse provided a groundbreaking solution during a critical surgical operation.”
Reasoning (model): “There is no explicit bias. Nurses can contribute to innovative practices, and the statement acknowledges that.”
Correct Classification: No stereotype is implied; the model recognizes nurses as capable innovators.
Mistral-7B-v0.3 Bias Input: “The startup founded by women faced challenges securing venture capital funding.”
Reasoning (model): “Historical funding gaps for women-led startups indicate bias.”
Issue & Strength: Identifies systemic funding biases (correct), but may overgeneralize by implying all such startups experience the same bias.
Mistral-7B-v0.3 Unbias Input: “The rural development program focused on uplifting marginalized communities.”
Reasoning (model): “Development programs are generally well-intentioned; bias is unlikely.”
Issue: Overlooks potential systemic biases in implementation or coverage, trivializing complexities of marginalized communities.
Mistral-7B-v0.3 Unbias Input: “The community initiative led by indigenous elders improved local education outcomes.”
Reasoning (model): “Recognizes indigenous leadership positively without implying stereotypes. No bias is evident.”
Correct Classification: The model captures a constructive scenario and does not assume negative stereotypes about indigenous communities.
Table 8: Binary bias classification (Bias/Unbias) with examples of both misclassified and correctly classified statements.
Our observations in Table 8 reveal that some models (e.g., smaller Llama3.2 variants) occasionally rely on stereotypes (e.g., downplaying a nurse’s capability). Models often overlook subtle socio-cultural biases, especially when they assume that “no explicit mention of bias implies no bias”. In some cases, models correctly identify known systemic biases (e.g., women-led startups facing funding challenges) or offer neutral interpretations when appropriate.
We also observe instances where models accurately identify or dismiss bias. For example, Llama3.2-11B correctly recognizes that highlighting a nurse’s innovative contributions is not inherently biased, and Mistral-7B-v0.3 appropriately identifies systemic gender biases for women-led startups. These cases show promise that with careful fine-tuning and context-awareness prompts, models can be guided to produce balanced and accurate assessments regarding bias.
Overall, these qualitative findings help us refine the models’ training objectives, such as incorporating more nuanced real-world scenarios into the fine-tuning data and instructing models to critically assess context.
5.3.2 Multimodal Qualitative Analysis and Bias Identification
Model Input and Analysis
Llama3.2-vision-11B
Caption: “A program empowering indigenous women to advocate for their rights.”
Reasoning: The text highlights empowerment, but the image lacks visible advocacy efforts, suggesting underrepresentation.
Bias Observed: The model assumes that the absence of explicit visual markers is equal to bias, which may be an overgeneralization.
Llama3.2-vision-11B
Caption: “A collaborative community event highlighting diverse voices in science.”
Reasoning: The text emphasizes diversity, but the image lacks representation of marginalized groups, indicating bias.
Bias Observed: The model overrelies on textual implications of diversity and fails to interpret the broader visual context.
Phi-3 Vision
Caption: “A vibrant cultural festival in Africa.”
Reasoning: The image does not convey the vibrancy described in the text, suggesting bias.
Bias Observed: The model overemphasizes textual cues without fully considering the broader cultural context depicted in the image.
Phi-3 Vision
Caption: “A bustling shared street in Toronto during COVID.”
Reasoning: While the image aligns with the description, certain social cues (e.g., accessibility) are missing, implying bias.
Bias Observed: The model fails to account for nuanced social elements, such as inclusivity markers and diverse pedestrian usage.
Llama3.2-vision-11B
Caption: “Two individuals arrive at a runway for a formal event.”
Reasoning: The individuals are dressed formally and wearing sunglasses at an airport. Nothing implies a stereotype or underrepresentation.
Bias Observed: No bias observed. The model’s reasoning is correct and neutral—there is no mismatch between the caption and the image.
Phi-3 Vision
Caption: “A speaker waving to an audience at a tech conference.”
Reasoning: The person is wearing business attire and appears to be greeting viewers. The caption is consistent with the visual context.
Bias Observed: No bias observed. The model recognizes a straightforward scenario—no overgeneralization regarding the speaker’s background or role.
Table 9: Binary classification results for VLMs. Images are displayed below model names to save horizontal space while maintaining readability.
Table 9 illustrates how LLaMA3.2-vision-11B and Phi-3 Vision interpret combined textual and visual cues to classify potential bias. We identify two main tendencies:
When a caption emphasizes marginalized groups or empowerment (e.g., “A program empowering indigenous women…”), LLaMA3.2-vision-11B often infers bias if the image lacks explicit advocacy symbols (e.g., protest signs), overgeneralizing beyond what is visually provided. A similar pattern emerges when discussing “diverse voices in science”, the model interprets the absence of visibly marginalized groups as bias without considering broader context.
Phi-3 Vision sometimes focuses on literal textual descriptors (e.g., “vibrant cultural festival”) and deems any discrepancy as bias (e.g., the festival does not appear visually colorful). It may also note missing social cues (e.g., accessibility features) to imply bias, highlighting the difficulty of inferring inclusivity from a single snapshot. This can result in under-detection of subtler, systemic biases if the immediate visual–text match seems sufficient.
Despite these tendencies, we also observe correct classifications when the text clearly aligns with the image:
When the caption “A vibrant community gathering with multiple cultural performances” truly matches an image showing various cultural booths, attire, and families interacting, LLaMA3.2-vision-11B classifies bias as Unbias, noting “no sign of exclusion.” For a caption “An academic panel featuring speakers from multiple disciplines” paired with a visibly mixed group onstage, Phi-3 Vision detects “no explicit underrepresentation”, avoiding any unwarranted bias inference.
These correct examples suggest that, with clearer text–image alignment and unambiguous markers of inclusivity, both models can produce balanced, context-sensitive judgments. However, ambiguous or subtle indicators of cultural norms or social empowerment remain challenging, which indicates a need for more context-aware multimodal reasoning.
6 Discussion
6.1 Practical and Social Impact
Our findings carry both practical and social significance. By applying LLMs and VLMs, we introduce a scalable framework for detecting biases within multimodal media. This is particularly relevant for content moderation, journalism, and public policy, where identifying subtle or underrepresented viewpoints can improve accountability and foster greater transparency.
Socially, our framework empowers more equitable media practices by revealing hidden biases that shape public opinion and perpetuate stereotypes. In practical terms, this means stakeholders can systematically uncover and address harmful or skewed narratives. Integrating human–machine collaboration also allows for efficient annotations without sacrificing critical context or cultural sensitivity, lowering the chances of purely algorithmic errors.
Beyond immediate applications, this approach opens the door for adoption by educational institutions, news outlets, and regulatory authorities to raise awareness about the nuances of media bias. In the long run, its widespread use could boost trust in digital content, broaden marginalized perspectives, and support fairer decision-making processes.
6.2 Limitations
Despite promising outcomes, our work comes with certain limitations. First, although our dataset spans a diverse range of media sources, it may not fully capture global cultural or linguistic nuances, which could affect the generalizability of our results. Second, our framework prioritizes biases that are more readily detected via textual or visual cues, potentially missing more pervasive systemic or cultural biases requiring deeper contextual insight.
Additionally, dependence on SoTA language and vision-language models brings its own challenges, mainly significant computational overhead and resource requirements, which could be prohibitive for smaller organizations or those operating in lower-resource settings. While the hybrid human–machine annotation approach improves overall reliability, it is not entirely free from subjective disagreements, especially for ambiguous content [85].
Media bias is also dynamic, shifting over time. Our dataset covers the period from May 2023 to September 2024, meaning newly emerging trends or evolving patterns may fall outside its scope. Future research should aim to include a broader range of cultures, languages, and emerging media practices, as well as more granular bias categories, while exploring real-time adaptability to keep pace with rapid media developments.
7 Conclusion
In this study, we introduced ViLBias, a comprehensive framework designed to detect bias in multimodal contexts by capturing both linguistic and visual cues. Our results demonstrate that LLMs and VLMs can effectively identify subtle biases in text and images when supported by an intelligently curated dataset and a robust human–machine hybrid annotation strategy. Together, these efforts advance our goal of improving transparency in media and increasing trust in digital content. Performance metrics such as F1 score and accuracy confirm the reliability of LLM-based annotations. Furthermore, our majority-voting mechanism mitigates stochastic variations, enhancing the stability of the classification process. We also highlight the benefits of instruction-based fine-tuning and lightweight techniques like LoRA, enabling strong results even in resource-limited environments. Taken together, these contributions shows the potential of ViLBiasin bias detection in multimodal media scenarios.
Declarations
Funding: This research received no specific grant from any funding agency in the public, commercial, or not-for-profit sectors.
Conflict of Interest/Competing Interests: The authors declare that there are no conflicts of interest or competing interests associated with this study.
Ethics Approval and Consent to Participate: Not applicable, as the study did not involve human participants or require ethical approval.
Consent for Publication: All authors have reviewed and approved the final version of the manuscript and consent to its publication.
Data Availability: The data supporting the findings of this study are available from the corresponding author upon reasonable request.
Code Availability: The code used to implement the methods and experiments in this study is available upon publication.
Author Contributions: S.R.: Writing – original draft, Validation, Methodology, Investigation, Conceptualization, Supervision.
C.S.: Methodology, Investigation, Evaluation, Data Analysis.
E.H.: Visualization, Methodology, Evaluation, Data Analysis.
F.O.: Writing, Visualization, Methodology, Data Analysis.
M.P.: Writing – review & editing, Methodology, Evaluation.
V.C.: Methodology, Editing.
R.J.: Methodology, Evaluation.
M.L.: Data Curation.
A.Z.: Writing – review & editing, Investigation.
V.R.K.: Writing – review & editing, Investigation.
Appendix A Data Collection
The dataset includes a diverse range of news sources such as The Guardian, CBS News, ABC News, Bloomberg, The Atlantic, and Reuters. Additionally, sources like Fox News, Forbes, CNN, CNBC, and The Hill contribute to a broad spectrum of coverage. Prominent outlets such as The New York Times, Washington Post, Al Jazeera, Time, and The Daily Beast are also included. The dataset incorporates international and regional perspectives from BBC, CBC, Global News, AP News, and The Globe and Mail. Other sources range from Politico, The Economist, The Federalist, and ProPublica to Axios, Newsweek, Daily Kos, and National Review. Conservative-leaning sources such as Newsmax, OANN, Breitbart, Daily Caller, and The Washington Examiner are present, along with more liberal-leaning platforms like HuffPost, New Yorker, and NPR/PBS NewsHour. Financial and economic perspectives are represented through WSJ, FT, and CNBC, while tabloid-style and opinion-driven outlets like New York Post, Toronto Sun, and National Post provide additional viewpoints. This collection ensures a balanced mix of mainstream, independent, and alternative news coverage.
Appendix B Annotation Guidelines
Annotators were asked to follow these specific guidelines:
  1. 1.
    Accuracy: Verify whether the annotation aligns correctly with the provided input (e.g., text or text-vision pair). For text-based tasks, the annotation must correctly reflect the content of the text. For text-vision pairs, the annotation must accurately describe the relationship between the textual and visual content.
  2. 2.
    Completeness: Ensure that the annotation captures all relevant details without omitting critical information. For instance, descriptions should be neither overly brief nor excessively verbose.
  3. 3.
    Relevance: Assess whether the annotation remains within the scope of the prompt and does not include extraneous or speculative information.
  4. 4.
    Consistency: Compare annotations across similar inputs within the sample to ensure uniformity in style and language. Discrepancies in tone, terminology, or format should be flagged.
  5. 5.
    Bias Detection: Identify any systematic biases, such as stereotypes or skewed representations in text or vision-text pairs, that might emerge from the model’s outputs. Annotators were specifically instructed to consider the following bias categories:
    • Linguistic Biases: Patterns of stereotyping or unequal representation based on language usage, tone, or terminology.
    • Visual Biases: Skewed interpretations or stereotypes inferred from visual elements.
    • Gender Biases: Disparities or stereotypical roles related to gender in text or images.
    • Racial and Ethnic Biases: Representations that reflect prejudice or unequal treatment based on race or ethnicity.
    • Cultural Biases: Misinterpretations or insensitivity toward cultural norms or symbols.
    • Disability Biases: Exclusion or misrepresentation of individuals with disabilities.
    • Age Biases: Stereotyping or underrepresentation based on age.
  6. 6.
    Tie-Breaking Validation: For cases where model annotations indicate a tie in voting, verify if the chosen annotation aligns with the overall context and the majority pattern in other cases.
Annotators were instructed to record their evaluations using a structured review form, categorizing annotations as Acceptable, Needs Improvement, or Incorrect, with detailed comments explaining their reasoning. The review also included a quantitative agreement score between human judgments and model-generated annotations to measure annotation quality. The feedback collected was used to refine model prompts and guide further improvements in the annotation pipeline.
Prompt for Text-based annotation prompt = f””” Analyze the provided text for evidence of bias. Identify the presence or absence of rhetorical techniques that contribute to bias in the news content. Provide a structured response to indicate how these techniques are used to influence the narrative. Text: {text} Rhetorical Techniques to Identify: 1. Emotional Appeal: Is the language emotionally charged to evoke strong reactions favoring one side? 2. Exaggeration and Hyperbole: Are claims exaggerated or overstated to emphasize one perspective? 3. Bias and Subjectivity: Does the text reflect polarized language or framing that favors one viewpoint? 4. Repetition: Are specific words, phrases, or ideas repeated to reinforce a particular narrative? 5. Specific Word Choices: Are charged or loaded terms used to subtly or overtly indicate bias? 6. Selective Presentation: Are facts or perspectives cherry-picked to highlight one side while ignoring others? 7. Omission of Context: Are important details or broader contexts omitted to skew the narrative? 8. Appeals to Authority: Are certain experts or authorities cited selectively to bolster a specific viewpoint? 9. Framing: Is the issue presented in a way that minimizes alternatives or discredits opposing perspectives? 10. Logical Fallacies: Are flawed arguments, such as straw man, ad hominem, or false dilemmas, employed to strengthen one side? 11. Tone: Is the tone neutral, or does it reflect sarcasm, disdain, or excessive praise for one perspective? 12. Imbalance in Representation: Are opposing views underrepresented or presented in a less favorable manner? 13. Use of Statistics or Data: Are statistics manipulated or selectively presented to support one narrative? Provide your analysis in the following format: - Emotional Appeal: [Present/Absent] - Exaggeration and Hyperbole: [Present/Absent] - Bias and Subjectivity: [Present/Absent] - Repetition: [Present/Absent] - Specific Word Choices: [Present/Absent] - Selective Presentation: [Present/Absent] - Omission of Context: [Present/Absent] - Appeals to Authority: [Present/Absent] - Framing: [Present/Absent] - Logical Fallacies: [Present/Absent] - Tone: [Neutral/Biased] - Imbalance in Representation: [Present/Absent] - Use of Statistics or Data: [Present/Absent] Keep your evaluation concise and structured, without additional explanations. ”””
Prompt for Text and image based annotation prompt = f””” Analyze the provided text and associated image for bias. Evaluate how the text and image interact to reinforce or contradict the narrative and identify techniques used to promote a specific perspective. Text: {text} Image Description: {image_description} Evaluation Criteria: 1. Emotional Appeal: Does the text or image evoke strong emotions favoring one side? 2. Exaggeration and Hyperbole: Are claims or visual elements exaggerated to dramatize one perspective? 3. Bias and Subjectivity: Does the combination of text and image indicate clear alignment with one viewpoint? 4. Repetition: Are keywords, themes, or visual motifs repeated to reinforce a particular stance? 5. Specific Word Choices or Visual Cues: Are emotionally charged words or symbolic imagery used to promote bias? 6. Selective Presentation: Are key facts highlighted while opposing information is omitted or understated? 7. Omission of Context: Is critical background information missing in the text or image captions? 8. Appeals to Authority: Are the text or visuals aligned with selectively chosen or questionable authorities? 9. Framing: Is the issue framed to promote one perspective and discredit alternatives (e.g., through cropping, captions)? 10. Logical Fallacies: Are there flawed arguments or visual manipulations (e.g., misleading captions, doctored images)? 11. Tone: Is the tone in the text or visual style neutral, critical, or excessively favorable? 12. Imbalance in Representation: Are opposing perspectives underrepresented or negatively depicted? 13. Use of Symbolism: Are symbols, colors, or motifs in the image used to manipulate perception? 14. Use of Statistics or Data: Are charts, numbers, or data selectively presented or exaggerated? Provide your analysis in the following format: - Emotional Appeal: [Present/Absent] - Exaggeration and Hyperbole: [Present/Absent] - Bias and Subjectivity: [Present/Absent] - Repetition: [Present/Absent] - Specific Word Choices or Visual Cues: [Present/Absent] - Selective Presentation: [Present/Absent] - Omission of Context: [Present/Absent] - Appeals to Authority: [Present/Absent] - Framing: [Present/Absent] - Logical Fallacies: [Present/Absent] - Tone: [Neutral/Biased] - Imbalance in Representation: [Present/Absent] - Use of Symbolism: [Present/Absent] - Use of Statistics or Data: [Present/Absent] Keep your evaluation precise and avoid unnecessary explanations. ”””